OpenStudy (anonymous):

What is the equation of a circle whose center is at the origin and whose radius is 9?

OpenStudy (anonymous):

\[(x-x_a)^2+(y-y_a)^2=r^2\\ (x_a,y_a)=(0,0)\\r=radius =9\]

OpenStudy (anonymous):

First line is the equation of a circle